### v2

# LVMark: Robust Watermark for latent video diffusion models

Abstract
Rapid advancements in generative models have made it possible to create hyper-realistic videos. As their applicability increases, their unauthorized use has raised significant concerns, leading to the growing demand for techniques to protect the ownership of the generative model itself. While existing watermarking methods effectively embed watermarks into image-generative models, they fail to account for temporal information, resulting in poor performance when applied to video-generative models. To address this issue, we introduce a novel watermarking method called LVMark, which embeds watermarks into video diffusion models. A key component of LVMark is a selective weight modulation strategy that efficiently embeds watermark messages into the video diffusion model while preserving the quality of the generated videos. To accurately decode messages in the presence of malicious attacks, we design a watermark decoder that leverages spatio-temporal information in the 3D wavelet domain through a cross-attention module. To the best of our knowledge, our approach is the first to highlight the potential of video-generative model watermarking as a valuable tool for enhancing the effectiveness of ownership protection in video-generative models. Project website: https://kuai-lab.github.io/lvmark2024/
1 Introduction
Video generative models have significantly enhanced their ability to create high-quality videos [54, 24, 3, 6, 44]. As these advancements increase the utility of video diffusion models, it is necessary to address the issue of model ownership infringement [12, 41, 18]. Since existing video watermarking methods embed the watermark into the video rather than the generative model [23, 39], they require costly post-processing for each video and cannot verify ownership if the generative model is stolen. As a way to establish ownership of generative models, embedding watermarks directly into diffusion models has recently collected much attention [18, 12]. Inspired by watermarking methods for image diffusion models, in this paper, we propose a method to embed watermarks into video diffusion models to establish model ownership.
Figure 1: Overview of video generation and ownership identification. Top: A binary message is embedded in the latent decoder to generate a watermarked video. Bottom: The watermark decoder extracts message from modified videos (e.g., H.264 compression, frame drops, cropping) to identify the model ownership.
To the best of our knowledge, there does not exist a watermarking method for watermarks into video-generative models, but several approaches that embed watermarks into image diffusion models have been proposed so far [12, 42]. The key challenge is how to embed a watermark with minimal degradation of generated images. The common idea is to fine-tune the latent decoder of the image generator using loss functions focused on image quality and watermark accuracy; for example, Stable Signature [12] employ LPIPS [52] to assess watermarked image quality and binary cross-entropy loss for the watermark message.
In the absence of watermarking methods for video diffusion models, applying the watermarking methods from image to video diffusion models is essential. However, watermarking methods for image diffusion models do not consider video-specific features, such as temporal consistency and robustness to video-specific attacks. When image diffusion models [29, 32, 34] are applied to video diffusion models, they fine-tune image-generative models to embed messages into each video frame using a 2D watermark decoder, which results in message embedding without considering inter-frame relationships. Moreover, these methods are vulnerable to video-specific attacks, such as H.264 compression, which is the widely used compression technology that easily removes messages embedded in the videos [23, 8, 37, 43, 20]. As these methods are trained to embed a single message into the model, they require a separate training process for each watermark message. Therefore, simply applying watermarking methods developed for image diffusion models is insufficient for video diffusion models.
To address these challenges, we introduce LVMark, a novel watermarking method specifically designed for video diffusion models. The goal of LVMark is to embed random binary messages into each frame of a video, enabling the generation of videos with different watermark messages after a single training process. To embed messages effectively while preserving video quality, we selectively modulate weight parameters in the latent decoder with minimal visual impact, enabling direct embedding of multiple random messages with single training [48]. Additionally, we design a robust watermark decoder that integrates the 3D wavelet domain which has spatio-temporal information with the RGB domain through a cross-attention module, which is shown to have the ability to effectively fuse two domains [16, 15]. To balance the trade-off between visual quality and message decoding accuracy, which is common in watermarking tasks, we propose a weighted patch loss that places more emphasis on patch areas where artifacts occur. Experimental results demonstrate that our method outperforms all the reasonable baselines. Since our method does not alter the structure of the diffusion model, it prevents anonymous users from bypassing the watermarking process. With its ability to identify AI-generated videos and verify ownership, LVMark protects ownership of the creations. Our contributions in this paper can be summarized as follows:
  • We propose a novel watermarking method that embeds watermarks into video diffusion models, to the best of our knowledge, for the first time.
  • By leveraging an importance-based weight modulation technique, our method embeds multiple messages into video diffusion models with a single training process while preserving video quality.
  • Our watermark decoder is constructed to efficiently leverage the 3D wavelet domain, capturing the spatio-temporal information necessary for robust message decoding against video-specific attacks, including H.264 compression and frame drops.
  • To balance visual quality with message decoding accuracy, we introduce a weighted patch loss that effectively reduces artifacts in localized areas.
2 Related Work
2.1 Video Diffusion Models
Diffusion models have become a powerful generative approach, refining noisy data progressively to produce high-quality outputs [36, 34, 27, 14, 46]. Latent Video Diffusion Models (LVDMs) [4] optimize spatio-temporal features within a compact latent space, reducing computational costs. OpenSora [54] and Latte [24] utilize the DiT to model long-range temporal dependencies [28]. In contrast, Stable video diffusion [3] and VideoCrafter [6] leverage U-Net to effectively capture both spatial and temporal dependencies. DynamiCrafter [44] enhances U-net based methods by modeling dynamic scenes and complex motion [50]. Both U-Net and DiT architectures play crucial roles in advancing video generation, highlighting the need for watermark methods that can be applied to both types of video diffusion models.
2.2 3D Discrete Wavelet Transform
3D discrete wavelet transform (3D DWT) is a classical method widely used to analyze temporal information of videos. 3D DWT operates along the time, height, and width of videos to decompose a low frequency subband and several high frequency subbands: $LLL$, $LLH$, $LHL$, $LHH$, $HLL$, $HLH$, $HHL$, and $HHH$ each with a size of $F/2\times H/2\times W/2$, where $L$ and $H$ denote low and high frequency, while $F$, $H$, and $W$ are frame, height, and width of video respectively. This operation analyzes the spatio-temporal information of videos. Existing video processing methods [51, 11, 49, 26, 47] utilize this property of 3D DWT to detect temporal information. Video digital watermarking methods [2, 22, 35, 17, 21] applied 3D DWT to the video for robustness of watermark. To the best of our knowledge, no research has yet explored watermarking video diffusion models using 3D DWT, underscoring the need for investigation to enhance robustness against common video modifications.
Figure 2: Training Pipeline. The training pipeline of our method is illustrated here. Top: We fine-tune the latent decoder to embed binary messages in generated videos and train the watermark decoder to retrieve messages from distorted videos. Bottom-left: We modulate layers of the latent decoder that minimally impact visual quality to embed random messages. Bottom-right: The watermark decoder combines the RGB video with low frequency subbands from a 3D wavelet transform using cross-attention to decode the binary message.
2.3 Diffusion Model Watermark
Digital watermarking plays a crucial role in protecting intellectual property, ensuring content authenticity, and enabling ownership tracking. Recently, digital watermarking technology has been applied to diffusion models. Watermarking methods for diffusion models [42, 53, 45] embed watermarks into the generated content from diffusion models without a post-processing step. However, since they mainly focus on the image-generative models, these methods perform poorly when applied to video diffusion models (Section 5.4), which highlights the need for research on watermarking in video-generative models.
3 Preliminaries
We utilize latent video diffusion models, which consist of a latent encoder $\mathcal{E}$, a latent decoder $\mathcal{D}$, and a diffusion model, which is typically based on one of two architectures: U-Net [33] or DiT [28]. The latent encoder $\mathcal{E}(v)$ maps the video $v\in\mathbb{R}^{d_{v}}$ to a latent representation $z\in\mathbb{R}^{d_{z}}$, where $d_{v}$ and $d_{z}$ denote the dimensions of video $v$ and latent vector $z$. In the forward process, noise is added to $z$, resulting in $z_{t}$, which is then denoised in the backward process to recover $z$, conditioned on inputs (e.g., text prompts). Finally, the latent decoder $\mathcal{D}$ converts $z$ back to the video $\hat{v}=\mathcal{D}(z)$. U-Net-based diffusion models focus on local features through downsampling and upsampling layers. In contrast, DiT-based models capture global dependencies and long-range correlations, making them effective for modeling complex temporal relationships. While image diffusion model watermarking methods [12, 41, 31] target U-Net-based models, we introduce a novel watermarking method applicable to both U-Net and DiT-based video diffusion models.
4 Methods
An overview of our method is shown in Figure 2. Our method fine-tunes the latent decoder $\mathcal{D}$ to generate video frames while embedding invisible binary messages into each frame.
4.1 Importance-based Weight Modulation
Our method aims to embed random messages into video diffusion models, enabling the generation of videos with different watermark messages. To achieve this, inspired by Yu et al. [48], we modulate the weight parameters of pre-trained latent decoder $\mathcal{D}$ based on importance of layers, as shown in bottom-left of Figure 2. For every iteration, we first transform the random message $m$ into the mapped message $m^{\prime}$ with the mapping network $\mathcal{M}(m):\mathbb{R}^{d_{m}}\rightarrow\mathbb{R}^{d_{m^{\prime}}}$, which is composed of two fully connected layers. Then, we modulate the $i$-th weight parameters of decoder $\mathcal{D}^{(i)}$ by performing element-wise multiplication between $\mathcal{D}^{(i)}$ and the mapped message, where the $i$-th fully connected layer $\mathcal{FC}^{(i)}(m^{\prime}):\mathbb{R}^{d_{m^{\prime}}}\rightarrow\mathbb{R}^{d_{\mathcal{D}}^{(i)}}$ is applied to match the dimension, where $d_{\mathcal{D}}^{(i)}$ is the input channel dimension of $i$-th weight parameters $\mathcal{D}^{(i)}$. Normalization is applied to ensure consistent scaling of the mapped message. The modulated weights in the $i$-th layer are defined as:
$\mathcal{D}^{(i)}_{mod}=\mathcal{D}^{(i)}\circ\big{(}\frac{\mathcal{FC}^{(i)}(\mathcal{M}(m))-\mu^{(i)}}{\sigma^{(i)}+\epsilon}\cdot\gamma^{(i)}+\beta^{(i)}\big{)},$ (1)
where $\gamma^{(i)}$ and $\beta^{(i)}$ are the $i$-th learnable parameters, while $\mu^{(i)}$ and $\sigma^{(i)}$ represent the mean and standard deviation of the message $\mathcal{FC}^{(i)}(\mathcal{M}(m))$. In our experiments, we find that modulating every layer of the decoder $\mathcal{D}$ leads to video quality degradation because it alters the parameters that have a significant impact on video generation. To minimize the undesirable impact on visual quality, we introduce importance-based weight modulation, which targets only the parameters with minimal effect on visual quality. To determine which layers to modulate, we first measure their importance. Specifically, before training, we sequentially add perturbations drawn from a normal distribution, $p\sim\mathcal{N}(0,1)$, to each layer, resulting in the perturbed weights $\mathcal{D}^{(i)}_{perturbed}=\mathcal{D}^{(i)}+p$. By evaluating the LPIPS [52] metric, which measures perceptual similarity, on videos generated with sequentially perturbed layers, our method identifies which layers have a greater impact on video quality. It modulates the least important 50% of the layers, which have less impact on the video quality.
4.2 Distortion Layer
Robustness against various video attacks is essential yet remains challenging for existing watermarking methods [23, 25]. To address this, we apply random distortions to the generated video in every iteration before extracting the message. We employ two types of distortions: spatial distortions and compression distortions. Spatial distortions refer to pixel-level distortions within each frame, such as cropping and blurring, which can be easily integrated. However, adapting to compression distortions, particularly H.264 compression, poses challenges during training due to their non-differentiable nature. To address this issue, we design an H.264 network inspired by DVMark [23] and train it to accurately replicate real H.264 compression, enhancing the robustness against H.264 compression.
4.3 Robust Video Watermark Decoder
As seen in the bottom-right of figure 2, we design a watermark decoder $\mathcal{W}$ that extracts messages by leveraging spatio-temporal information through the fusion of two domains: the generated RGB video $\hat{v}$ and its 3D wavelet transformed subband $\hat{v}_{freq}$. We use only the low frequency subband $LLL$ among the 3D wavelet subbands to enhance robustness against video compression, as various compression methods discard high-frequency features [19].
The watermark decoder is composed of two primary components: a feature extraction module and a feature fusion module. Before combining RGB domain and frequency domain, we first extract each domain’s features independently. Using six 3D convolutional blocks, we extract wavelet features $F_{freq}\in\mathbb{R}^{B\times 2048\times F\times H\times W}$ from the low frequency subband $LLL$, where $B$, $F$, $H$, and $W$ denote batch, frame, height, and width, respectively. On the other hand, we employ the ResNet50 model to extract spatial features $F_{rgb}\in\mathbb{R}^{B\times 2048\times F\times H\times W}$ from the RGB video $\hat{v}$.
To efficiently fuse these two features, we employ temporal and spatial cross-attention mechanisms. The temporal cross-attention measures the attention score across frames, while the spatial cross-attention measures the attention score across height and width. The temporal and spatial cross-attention modules share the same architecture, each designed with 8 multi-heads. The entire feature fusion module consists of temporal cross-attention, spatial cross-attention, and a position-wise feed-forward network, with a depth of 2. We set the wavelet features as the query $Q=\text{FC}_{q}(F_{freq})$, while using the spatial features as the key $K=\text{FC}_{k}(F_{rgb})$ and the value $V=\text{FC}_{v}(F_{rgb})$, where $\text{FC}_{q}$ , $\text{FC}_{k}$ and $\text{FC}_{v}$ denote a fully connected layer for query, key, and value, respectively. This allows the spatial feature $F_{rgb}$ to capture information from the wavelet features $F_{freq}$ by leveraging the attention scores computed between them. The cross-attention module $CA$ is defined as:
$CA(F_{freq},F_{rgb})=\text{softmax}(\frac{Q\cdot K^{T}}{\sqrt{d_{k}}})\cdot V.$ (2)
The output of each cross-attention module is added to the spatial feature $F_{rgb}$ for residual connection. With this watermark decoder, we obtain the message hidden within the generated video.
4.4 Training Objectives
We design our objective function to balance the trade-off between visual quality and watermark decoding accuracy. We employ binary cross-entropy loss between the random message $m$ and the extracted message $\hat{m}$ from the generated video to ensure accurate watermark decoding:
$\mathcal{L}_{message}=-\frac{1}{N}\sum_{i=1}^{N}$ $m_{i}\cdot\log(\sigma(\hat{m}_{i}))$ (3)
$+(1-m_{i})\cdot\log(1-\sigma(\hat{m}_{i})),$
where $N$ is the total number of message bits, while $\sigma(\cdot)$ denotes the sigmoid function. In addition, to enhance the visual quality of the generated video, we choose the Watson-VGG perceptual loss [10], which is designed to avoid blurring effects and visual artifacts of generated images:
$\mathcal{L}_{lpips}=\text{VGG}(\mathcal{D}_{mod}({\mathcal{M}}(m),{\mathcal{E}}(v)),v),$ (4)
where $\mathcal{D}_{mod}$, $\mathcal{M}$, and $\mathcal{E}$ denote the modulated latent decoder, mapping network, and latent encoder, respectively, while $m$ and $v$ are the random message and input video.
Furthermore, we propose a weighted patch loss to address localized artifacts in the generated video. We first compute the mean absolute error (MAE) for video patches, resulting in a loss score map normalized between 0 and 1 using the softmax function. By multiplying the patch MAE by the score map, we create a weighted patch loss that places more emphasis on areas with higher MAE, where artifacts are likely to occur. The weighted patch loss is as follows:
$\mathcal{L}_{patch}$ $=\frac{1}{P}\sum_{i=1}^{P}|\hat{v}_{i}-v_{i}|\cdot\text{softmax}(|\hat{v}_{i}-v_{i}|),$ (5)
where $P$ denotes the number of patches. The final objective can be formulated as:
$\mathcal{L}=$ $\lambda_{message}\mathcal{L}_{message}$ (6)
$+\lambda_{lpips}\mathcal{L}_{lpips}+\lambda_{patch}\mathcal{L}_{patch},$
where $\lambda_{message}$, $\lambda_{lpips}$, and $\lambda_{patch}$ are hyperparameters that balance three loss functions.
Figure 3: Qualitative results with baselines. We show the visual quality of generated videos with baseline watermarking methods. The first row shows the crop of each image, while the second and third rows show the video frame itself and the difference map($\times$10) between the original and each method.
Open Sora DynamiCrafter
Bit Method Bits Acc.(%) $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ tLP $\downarrow$ FVD $\downarrow$ Bits Acc.(%) $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ tLP $\downarrow$ FVD $\downarrow$
32 HiDDeN [55] 90.83 30.20 0.901 0.265 5.493 2707.9 90.48 30.19 0.887 0.187 5.410 1637.7
Blind [25] 92.05 28.88 0.888 0.251 5.615 1278.9 87.56 26.44 0.852 0.239 5.348 1390.2
Stable Signature [12] 86.32 21.71 0.681 0.285 3.307 494.8 87.65 21.39 0.541 0.162 0.511 249.2
WOUAF [18] 87.28 23.62 0.838 0.166 0.856 193.7 90.62 22.38 0.839 0.125 0.763 469.4
Ours 94.44 31.15 0.919 0.105 0.219 65.73 96.00 31.50 0.892 0.075 0.173 85.49
48 HiDDeN [55] 85.62 29.85 0.888 0.269 6.802 2731.9 86.96 29.83 0.876 0.188 5.843 1716.5
Blind [25] 87.41 27.79 0.872 0.204 4.845 1365.8 86.14 25.23 0.848 0.259 5.480 1480.1
Stable Signature [12] 78.99 20.57 0.614 0.297 4.019 688.5 81.48 20.82 0.530 0.175 0.603 262.8
WOUAF [18] 83.05 23.47 0.828 0.168 1.053 196.2 89.63 22.23 0.811 0.130 0.789 523.2
Ours 90.42 30.32 0.910 0.117 0.287 117.5 93.67 30.78 0.889 0.100 0.240 155.0
Table 1: Quantitative results. We present the quantitative results for Open-Sora [54] and DynamiCrafter [44] with various watermarking methods: HiDDeN [55], Blind [25], Stable Signature [12], WOUAF [18] and LVMark, applied. The evaluation includes image metrics: PSNR, SSIM, LPIPS, video metrics: tLP [9], FVD [38], and bit accuracy on 32 and 48-bit messages.
5 Experiments
5.1 Dataset
Video dataset. We train our method on Panda-70M dataset [7], a large-scale collection of video frames recognized for its diversity across various content types. We randomly download 10K videos from this dataset. Following the approach of Stable Video Diffusion [3], we sample 8 frames from each video, with each frame resized to a resolution of $256\times 256$ for training.
Prompt dataset. We evaluate text-to-video generation on the Vidprom dataset [40], using prompts generated by GPT-4 [1]. To evaluate our method, we randomly sample 100 prompts from the dataset, which do not contain personally identifiable information or offensive content.
Bit Accuracy(%) $\uparrow$
None
Crop
(50%)
Rotation
($+\pi/6$)
Gaussian Blur
(std = 1.0)
JPEG Comp
(70%)
Frame Swap
(p=0.5)
Frame Avg.
(n=6)
Frame Drop
(p=0.5)
H.264
(crf=21)
Combine
Model Distortion
(p=0.2)
Open Sora HiDDeN [55] 90.83 90.71 62.26 75.06 71.18 90.83 90.77 90.88 61.70 61.72 -
Blind [25] 92.05 87.91 86.30 68.56 62.92 86.63 55.75 89.50 84.70 82.97 -
Stable Signature [12] 86.32 78.42 54.26 84.15 76.65 86.32 77.43 76.72 81.31 51.82 77.58
WOUAF [18] 87.28 79.79 78.62 86.94 85.66 87.28 81.91 82.19 81.24 76.21 75.33
Ours 94.44 91.85 87.10 93.21 87.94 94.44 93.88 94.40 90.21 83.97 89.69
Dynami Crafter HiDDeN [55] 90.48 90.45 50.32 78.01 82.24 90.48 88.47 90.44 62.15 61.97 -
Blind [25] 87.56 83.74 86.26 67.29 59.53 84.38 57.88 84.57 83.45 85.49 -
Stable Signature [12] 87.65 84.15 50.39 82.20 77.44 87.65 83.97 85.97 83.94 51.62 80.98
WOUAF [18] 90.62 90.59 89.03 88.41 87.99 90.62 88.61 90.50 89.70 88.42 78.36
Ours 96.00 93.42 93.02 95.18 94.04 96.00 95.48 94.95 95.05 91.91 90.23
Table 2: Robustness comparisons. We present bit accuracy results of watermarking methods under various attacks, including image distortions, video distortions, combined distortions, and model distortions.
5.2 Implementation Details
Video diffusion models are generally categorized into two types: DiT [28] and U-Net [33] architectures. To evaluate our method on DiT architecture, we employ Open-Sora [54], an open-source text-to-video-generative model. For the U-Net architecture, we utilize DynamiCrafter [44], which is the notable superior in video quality for text-image-to-video. We follow the same approach as DynamiCrafter by collecting images. Notably, these images are exclusively utilized in DynamiCrafter to generate videos for our evaluation process. For training, we only utilize the latent decoder of these models. Our method is trained on 10K videos (8 frames, $256\times 256$ resolution) from Panda-70M dataset [7]. We modulate 50 % of the latent decoder layers to embed the message. For optimization, we set $\lambda_{message}=0.8$, $\lambda_{lpips}=0.7$, $\lambda_{patch}=10$ in Eq. 6. Furthermore, we employ 100 random messages.
5.3 Evaluation
[…]
Baseline. To the best of our knowledge, there is no watermarking method for the video diffusion models. Therefore, we compare with four strategies for a fair comparison: 1) HiDDeN [55], is the first deep learning image watermarking method. We embed the same message into each video frame. 2) Blind [25], is the traditional video watermarking method, using 3D Discrete Wavelet Transform. We follow the implementation in Blind [25]. 3) Stable Signature [12], is the watermarking method for the text-to-image diffusion model. 4) WOUAF [18], is the state-of-the-art watermarking method for the text-to-image diffusion model. For both Stable Signature and WOUAF, we train Open-Sora [54] and DynamiCrafter [44] using the same training set as our method. Additionally, we decode the message from each video frame for HiDDeN, Stable Signature, and WOUAF.
5.4 Experimental results
Video Quality and Bit Accuracy. There is a trade-off between video quality and bit accuracy, making it a key challenge in digital watermarking to improve both simultaneously. We first compare the generated video quality and bit accuracy of our method against other approaches, as shown in Figure 3 and Table 1. We observe that Stable Signature [12] and WOUAF [18] fail to generate high-quality videos with an invisible message when extended to video generation. HiDDeN [55] does not fully address temporal consistency, as it independently embeds the invisible message into each frame of the generated video. Though Blind [25] is a traditional video watermarking method, it fails to ensure temporal consistency on synthetic videos, as demonstrated by tLP and FVD evaluations, highlighting its limitations in handling generated videos. In contrast, as shown in Figure 3, and Table 1, our method preserves video quality, including temporal consistency, while achieving the highest bit accuracy. To illustrate the differences between the original and watermarked video, we present the normalized pixel intensity difference in Figure 4, following previous watermarking visualization methods [5, 30, 13]. Our method shows the smallest difference, demonstrating superior visual quality (more details in Section 5.5).
Figure 4: Visualization of pixel intensity difference. We show the normalized pixel intensity difference between original and watermarked videos. The red line depicts our results. A lower value on the y-axis indicates greater similarity to the original video.
Video Quality and Capacity. In this section, we analyze video quality and bit accuracy based on the length of the embedded message. We compare our method with others for 32-bit and 48-bit messages, using 100 different messages. As shown in Table 1, all methods exhibit a decrease in bit accuracy as the message length increases. We note that our method achieves the highest bit accuracy compared to others, with over 90% bit accuracy at 48 bits. Furthermore, while all methods experience a noticeable drop in video quality as message length increases, our approach consistently shows a smaller decline across all metrics than other state-of-the-art methods. Our method also allows for embedding more messages in videos of comparable quality, outperforming others in both bit accuracy and video quality.
Robustness for Video Distortion. One of the priorities in digital watermarking is robustness against various distortions. In this section, we compare the robustness of our method with other approaches to demonstrate its effectiveness. As shown in Table 2, HiDDeN [55] is vulnerable to H.264 compression, while Blind [25] exhibits weaknesses against frame averaging, Gaussian blur, and JPEG compression. Additionally, Stable Signature [12] and WOUAF [18] fail to preserve the message when videos generated by the DiT model undergo rotation. In contrast, our method preserves the message across all attack metrics, thanks to the watermark decoder using 3D wavelet features.
Notably, it demonstrates exceptional robustness against video compression, especially H.264. We note that since all methods except Blind [25] are applied independently per-frame, it is naturally robust against frame drop.
Robustness for Model Distortion. We explore the scenario in which a malicious user, recognizing the existence of embedded messages in videos generated by the latent decoder, distorts the latent decoder to remove the embedded messages. We adopt a strategy that involves adding Gaussian noise ($\sigma=0.05$) to the parameters of the latent decoder. We randomly select 20 % parameters of the latent decoder. Table 1 shows that our method ensures the presence of messages even when the latent decoder is distorted.
5.5 Ablation study
For the ablation study, we use the same configuration as described in Section 5.2, with 32-bit messages.
Figure 5: The impact of weight modulation rate. Each point represents the metrics obtained by training with a specific modulation rate. We experiment with modulation rates of 0%, 25%, 50%, 75%, and 100%.
DWT Type Bits Acc. $\uparrow$ PSNR $\uparrow$ tLP $\downarrow$ FVD $\downarrow$
Open Sora 2D DWT 92.01 30.92 0.450 122.9
3D DWT (Ours) 94.44 31.15 0.219 65.73
Dynami Crafter 2D DWT 91.71 28.73 0.383 158.3
3D DWT (Ours) 96.00 31.50 0.173 85.49
Table 3: Comparison of different DWT types. We evaluate the metrics of 2D DWT and 3D DWT.
Bit Accuracy(%)    $\uparrow$
Domain Type PSNR $\uparrow$ None H264(crf21)
Open Sora RGB only 30.64 89.85 87.83
High Freq only 31.91 50.45 50.33
Low Freq only 33.30 50.13 50.12
High Freq + RGB 28.97 88.72 86.14
Low Freq + RGB (Ours) 31.15 94.44 90.21
Dynami Crafter RGB only 28.62 91.82 88.10
High Freq only 31.16 50.38 56.25
Low Freq only 29.79 51.04 50.19
High Freq + RGB 28.07 90.10 88.95
Low Freq + RGB (Ours) 31.50 96.00 95.05
Table 4: Comparison of different domain types. We evaluate the metrics for each domain type.
Weight modulation for watermarking As mentioned in Section 4.1, we selectively modulate the weights of the latent decoder to embed messages. The modulation rate is an important factor, as there exists trade-off. As shown in Figure 5, we generate videos with $k$ ranging from 0 to 100. As $k$ increases, bit accuracy improves while video quality declines, illustrating a trade-off. We find an optimal balance between bit accuracy and video quality when $k$ is set to 50.
Watermark Decoder with 3D DWT. Temporal consistency is a crucial factor in video watermarking and generation tasks. To address this, we employ the 3D Wavelet Transform (3D DWT), which captures spatio-temporal information and converts videos into wavelet signals. Table 3 shows the comparison between 3D DWT and 2D Discrete Wavelet Transform (2D DWT), which utilizes only spatial information. As shown in Table 3, our watermark decoder utilizing 3D DWT achieves the best temporal consistency, bit accuracy, and video quality.
Watermark Decoder with Low Frequency. As mentioned in Section 4.3, we initially utilize only the low frequency subband ($LLL$) among the 3D wavelet subbands. While this achieves high PSNR values, relying solely on the frequency domain limits bit accuracy, as shown in Table 4. By combining RGB features with the low frequency subband ($LLL$), we significantly enhance bit accuracy without compromising PSNR. This approach is particularly robust for decoding messages, even in H.264 compressed videos.
Figure 6: Visual impact of weighted patch loss. We visualize a video frame trained with and without weighted patch loss. To enhance clarity, we crop the local regions containing artifacts.
Open Sora DynamiCrafter
Bits Acc. $\uparrow$ LPIPS $\downarrow$ Bits Acc. $\uparrow$ LPIPS $\downarrow$
W/o Patch Loss 95.96 0.151 96.41 0.115
W/ Patch Loss (Ours) 94.44 0.105 96.00 0.075
Table 5: Ablation on weighted patch loss. We evaluate the impact of weighted patch loss.
Impact of Weighted Patch Loss on the video. As depicted in Section. 4.4, artifacts appear in localized areas of the generated videos. To address this issue, we design a loss function that focuses on the regions where artifacts occur. Without the patch loss, although we achieve higher bit accuracy, the visual quality significantly degrades due to the presence of local artifacts, especially in the DiT model: Open-sora, which relatively focuses less on the local area by the nature of transformer architectures. Table 5 and Figure 6 demonstrate the effectiveness of the weighted patch loss in eliminating artifacts from local regions.
6 Conclusion
We proposed LVMark, a first watermarking method for video diffusion models.
Our method captures spatio-temporal information by effectively fusing the 3D wavelet domain and the RGB spatial domain for accurate message extraction from generated videos, even in the case of video modifications. Additionally, our method generates high-quality watermarked videos by selectively modulating latent decoder weights according to the importance of each layer. By balancing the trade-off between video quality and bit accuracy with tailored objective functions, including weighted patch loss, LVMark successfully embedded watermarks into video diffusion models. Experimental results showed that LVMark outperforms all the reasonable baseline methods.
Limitations. Our method demonstrates outstanding performance in the digital watermarking of video diffusion models. However, training the diffusion model to embed the watermark requires substantial memory due to the large size of the video diffusion models and the videos themselves, with memory usage reaching approximately 25GB. Future work could investigate memory-efficient watermarking methods for video diffusion models.