### v3

# LVMark: Robust Watermark for Latent Video Diffusion Models

Abstract
Rapid advancements in video diffusion models have enabled the creation of realistic videos, raising concerns about unauthorized use and driving the demand for techniques to protect model ownership. Existing watermarking methods, while effective for image diffusion models, do not account for temporal consistency, leading to degraded video quality and reduced robustness against video distortions. To address this issue, we introduce LVMark, a novel watermarking method for video diffusion models. We propose a new watermark decoder tailored for generated videos by learning the consistency between adjacent frames. It ensures accurate message decoding, even under malicious attacks, by combining the low-frequency components of the 3D wavelet domain with the RGB features of the video. Additionally, our approach minimizes video quality degradation by embedding watermark messages in layers with minimal impact on visual appearance using an importance-based weight modulation strategy. We optimize both the watermark decoder and the latent decoder of diffusion model, effectively balancing the trade-off between visual quality and bit accuracy. Our experiments show that our method embeds invisible watermarks into video diffusion models, ensuring robust decoding accuracy with 512-bit capacity, even under video distortions.
1 Introduction
Video diffusion models have made significant progress in generating high-quality videos [45, 22, 3, 5, 37]. With these advancements, the need for methods to establish ownership of both the diffusion model and its creations has emerged. This issue can be addressed by embedding watermarks into the video diffusion model and enabling tracking.
Figure 1: Overview of video generation and ownership identification. Top: An authorized user generates a watermarked video. Bottom: Despite distortions applied to the video or model, the watermark decoder can reliably identify the video’s owner.
Existing watermarking methods applicable to video generation can be categorized into three types. The first approach uses post-processing techniques [46, 30, 23, 44]. While these methods enable accurate bit extraction, they can be easily bypassed if the diffusion model is stolen, as the watermark information is not part of the model. The second approach embeds watermarks into the latent noise of the diffusion model [35, 12]. This approach modifies the latent noise to generate altered content and requires the exact inverse process of the diffusion model for message extraction. Moreover, user-controlled embedding allows fingerprints to be excluded if the model is stolen [16]. The third approach fine-tunes a pre-trained diffusion model to embed the watermark [10, 16], ensuring watermark insertion cannot be bypassed while preserving the original content.
However, fine-tuning-based methods have only been explored for image diffusion models, not for videos. While 2D watermark decoders can be applied to individual frames, following the strategy of image diffusion watermarking methods, they fail to preserve temporal consistency and leverage the large data volume of videos for high-capacity watermarking. This underscores the need for a video-specific watermark decoder for video diffusion models.
To the best of our knowledge, this paper proposes the first fine-tuning-based watermarking method for video diffusion models, LVMark. We propose a new watermark decoder tailored for video diffusion models, focusing on temporal consistency. Extracting watermarks from the RGB domain forces the model to embed them in visible color regions, which degrades image quality and makes the watermark vulnerable to distortions. To address this, we integrate the 3D wavelet and RGB domains, embedding watermarks in the invisible frequency domain, with the cross-attention module [14, 13] enabling robust decoding. Additionally, to improve compression robustness, we use the low-frequency subband of the 3D wavelet domain, as video compression methods like H.264 tend to remove high-frequency regions. We optimize the watermark decoder alongside the latent decoder of the diffusion model by iteratively embedding and extracting random bits, ensuring resilient bit extraction despite weight variations.
An effective method to embed random watermarks in the diffusion model is weight modulation [16], which directly embeds watermarks into the model’s weight parameters. Despite the strong watermark embedding capability of this strategy, our experiments reveal that modulating all layers of the latent decoder degrades visual quality. To preserve video quality, we selectively modulate the latent decoder’s weight parameters with minimal visual impact. We evaluate the impact of each layer on video quality by injecting Gaussian noise into individual layers and measuring the LPIPS score [43]. To further enhance both visual quality and message decoding accuracy, we propose a weighted patch loss that focuses on regions where artifacts are more likely to occur, effectively balancing these two aspects.
Experimental results demonstrate that our method outperforms all reasonable baselines, particularly in terms of video consistency as well as robustness against attacks. Specifically, embedding a 512-bit watermark into DynamiCrafter [37] achieves a Fréchet Video Distance of 68.94, proving superior temporal consistency over state-of-the-art (SOTA) methods. After applying H.264 compression, crop, and frame drop attacks, our method achieves a bit accuracy of 99.16%, outperforming the SOTA method by 37.88%. We can establish model ownership even in cases of unauthorized use, as depicted in Figure 1, thanks to our method’s ability to embed high-capacity, imperceptible watermarks into video diffusion models.
Our contributions can be summarized as follows:
  • We introduce a robust watermarking method for video diffusion models.
  • We propose a watermark decoder that leverages low-frequency 3D wavelet components for robust spatio-temporal message decoding, ensuring resilience to video-specific attacks like H.264 compression and frame drops.
  • By leveraging an importance-based weight modulation technique, our method embeds multiple messages into video diffusion models with a single training process while preserving video quality.
  • To balance visual quality and message decoding accuracy, we introduce a weighted patch loss that effectively reduces localized artifacts.
2 Related Work
2.1 Video Diffusion Models
Diffusion models have become a powerful generative approach, refining noisy data progressively to produce high-quality outputs [31, 28, 25, 11, 38]. Latent Video Diffusion Models (LVDMs) [4] optimize spatio-temporal features within a latent space, reducing computational costs. OpenSora [45] and Latte [22] utilize the DiT [26] to model long-range temporal dependencies. In contrast, Stable video diffusion [3] and VideoCrafter [5] leverage U-Net [27] to effectively capture both spatial and temporal dependencies. DynamiCrafter [37] enhances U-Net by modeling dynamic scenes and complex motion [41]. Both U-Net and DiT architectures play crucial roles in advancing video generation, highlighting the need for watermark methods that can be applied to both types of video diffusion models.
2.2 Diffusion Model Watermark
Digital watermarking plays a crucial role in protecting intellectual property, ensuring content authenticity, and enabling ownership tracking. Watermarking methods for diffusion models [36, 16, 10] embed watermarks into the generated content from diffusion models without a post-processing step. However, since they mainly focus on image diffusion models, these methods perform poorly when applied to video diffusion models (Section 4.4). Although VideoShield [12] is the first to propose watermarking in video diffusion models, it alters the generated video by embedding the watermark in latent noise and requires the inverse process of the diffusion model for message decoding. These limitations highlight the need for watermarking techniques that do not alter the generated image and allow for a simple decoding process.
Figure 2: Training pipeline. The training pipeline of our method is illustrated here. Top: We fine-tune the latent decoder to embed binary messages in generated videos and train the watermark decoder to retrieve messages from distorted videos. Bottom-left: We modulate layers of the latent decoder that minimally impact visual quality to embed random messages. Bottom-right: The watermark decoder combines the RGB video with low-frequency subbands from a 3D wavelet transform using cross-attention to decode the binary message.
2.3 3D Discrete Wavelet Transform
3D discrete wavelet transform (3D DWT) is a classical method widely used to analyze temporal information of videos. 3D DWT operates along the time, height, and width of videos to decompose a low-frequency subband and several high-frequency subbands: $LLL$, $LLH$, $LHL$, $LHH$, $HLL$, $HLH$, $HHL$, and $HHH$ each with a size of $F/2\times H/2\times W/2$, where $L$ and $H$ denote low and high-frequency, while $F$, $H$, and $W$ are frame, height, and width of video respectively. Existing video processing methods [42, 9, 40, 24, 39] leverage the properties of 3D DWT to capture temporal information, while video digital watermarking techniques [2, 19, 29, 15, 18] utilize 3D DWT to enhance the robustness of the watermark. To the best of our knowledge, no research has yet explored watermarking video diffusion models using 3D DWT, underscoring the need for investigation to enhance robustness against common video modifications.
3 Methods
An overview of our method is shown in Figure 2. We introduce a novel watermark decoder that incorporates the video’s temporal consistency (Section 3.1). We selectively modulate the weights of the latent decoder $\mathcal{D}$ to embed imperceptible random watermarks into the model (Section 3.2). The distortion layer ensures robust watermark decoding against distortions like crop, rotation, and H.264 compression (Section 3.3). By iteratively optimizing the watermark and latent decoders with random watermark messages, our method embeds an imperceptible watermark while ensuring robust extraction (Section 3.4).
Preliminaries. We utilize latent video diffusion models, which consist of a latent encoder $\mathcal{E}$, a latent decoder $\mathcal{D}$, and a diffusion model, which is typically based on one of two architectures: U-Net [27] or DiT [26]. The latent encoder $\mathcal{E}(v)$ maps the video $v\in\mathbb{R}^{d_{v}}$ to a latent representation $z\in\mathbb{R}^{d_{z}}$, where $d_{v}$ and $d_{z}$ denote the dimensions of video $v$ and latent vector $z$. In the forward process, noise is added to $z$, resulting in $z_{t}$, which is then denoised in the backward process to recover $z$, conditioned on inputs (e.g., text prompts). Finally, the latent decoder $\mathcal{D}$ converts $z$ back to the video $\hat{v}=\mathcal{D}(z)$. U-Net-based diffusion models focus on local features through downsampling and upsampling layers. In contrast, DiT-based models capture global dependencies and long-range correlations, making them effective for modeling complex temporal relationships. While image diffusion model watermarking methods [10, 35] target U-Net-based models, we introduce a novel watermarking method applicable to both U-Net and DiT-based video diffusion models.
3.1 Robust Video Watermark Decoder
We propose a new watermark decoder $\mathcal{W}$ that extracts messages from generated videos by leveraging spatio-temporal information through the fusion of two domains: the generated RGB video $\hat{v}$ and its 3D wavelet transformed subband $\hat{v}_{freq}$, as depicted in the bottom-right of Figure 2. We use only the low-frequency subband $LLL$ among the 3D wavelet subbands to enhance robustness against video compression, as various compression methods discard high-frequency features [17].
The watermark decoder is composed of two primary components: a feature extraction module and a feature fusion module. Before combining the RGB domain and frequency domain, we first extract each domain’s features independently. Using six 3D convolutional blocks, we extract wavelet features $F_{freq}\in\mathbb{R}^{B\times 2048\times F\times H\times W}$ from the low-frequency subband $LLL$, where $B$, $F$, $H$, and $W$ denote batch, frame, height, and width, respectively. On the other hand, we employ the ResNet50 model to extract spatial features $F_{rgb}\in\mathbb{R}^{B\times 2048\times F\times H\times W}$ from the RGB video $\hat{v}$.
To efficiently fuse these two features, we employ temporal and spatial cross-attention mechanisms. The temporal cross-attention measures the attention score across frames, while the spatial cross-attention measures the attention score across height and width. The temporal and spatial cross-attention modules share the same architecture, each designed with 8 multi-heads. The entire feature fusion module consists of temporal cross-attention, spatial cross-attention, and a position-wise feed-forward network, with a depth of 2. We set the wavelet features as the query $Q=\text{FC}_{q}(F_{freq})$, while using the spatial features as the key $K=\text{FC}_{k}(F_{rgb})$ and the value $V=\text{FC}_{v}(F_{rgb})$, where $\text{FC}_{q}$, $\text{FC}_{k}$ and $\text{FC}_{v}$ denote a fully connected layer for query, key, and value, respectively. This allows the spatial feature $F_{rgb}$ to capture information from the wavelet features $F_{freq}$ by leveraging the attention scores computed between them. The cross-attention module $CA$ is defined as:
$CA(F_{freq},F_{rgb})=\text{softmax}\left(\frac{Q\cdot K^{T}}{\sqrt{d_{k}}}\right)\cdot V.$ (1)
The output of each cross-attention module is added to the spatial feature $F_{rgb}$ for residual connection. With this watermark decoder, we obtain the message hidden within the generated video.
3.2 Importance-based Weight Modulation
To embed random messages into the diffusion model, we propose an importance-based weight modulation module inspired by WOUAF [16], which modulates the weight parameters of the pre-trained latent decoder $\mathcal{D}$. Although WOUAF successfully embeds messages into the model, our experiments show that modulating all layers degrades visual quality by altering parameters crucial for video generation. To mitigate this, we propose a novel approach that modulates only a subset of layers with lower importance, as shown in bottom-left of Figure 2. As a pre-processing step, we quantify layer importance by sequentially perturbing each layer with noise sampled from a Gaussian distribution, $p\sim\mathcal{N}(0,0.04)$, resulting in perturbed weights: $\mathcal{D}^{(i)}_{perturbed}=\mathcal{D}^{(i)}+p$. By evaluating the LPIPS [43] metric, which measures perceptual similarity, on videos generated with sequentially perturbed layers, our method identifies which layers have a greater impact on video quality.
In the training phase, we transform the random message $m$ into the mapped message $m^{\prime}$ with the mapping network $\mathcal{M}(m):\mathbb{R}^{d_{m}}\rightarrow\mathbb{R}^{d_{m^{\prime}}}$, which is composed of two fully connected layers. Then, we modulate the $i$-th selected layer $\mathcal{D}^{(i)}$ by performing element-wise multiplication between $\mathcal{D}^{(i)}$ and the mapped message, where the $i$-th fully connected layer $\mathcal{FC}^{(i)}(m^{\prime}):\mathbb{R}^{d_{m^{\prime}}}\rightarrow\mathbb{R}^{d_{\mathcal{D}}^{(i)}}$ is applied to match the dimension, where $d_{\mathcal{D}}^{(i)}$ is the input channel dimension of $i$-th layer’s weight parameters $\mathcal{D}^{(i)}$. Normalization is applied to ensure consistent scaling of the mapped message. The modulated weights in the $i$-th layer are defined as:
$\mathcal{D}^{(i)}_{mod}=\mathcal{D}^{(i)}\circ\left(\tfrac{\mathcal{FC}^{(i)}(\mathcal{M}(m))-\mu^{(i)}}{\sigma^{(i)}+\epsilon}\cdot\gamma^{(i)}+\beta^{(i)}\right),$ (2)
where $\gamma^{(i)}$ and $\beta^{(i)}$ are the $i$-th learnable parameters, while $\mu^{(i)}$ and $\sigma^{(i)}$ represent the mean and standard deviation of the message $\mathcal{FC}^{(i)}(\mathcal{M}(m))$.
3.3 Distortion Layer
Robustness against various video attacks is essential yet remains challenging for existing watermarking methods [21, 23]. To address this, we apply random distortions to the generated video in every iteration. We employ two types of distortions: spatial distortions and compression distortions. Spatial distortions refer to pixel-level distortions within each frame, such as cropping and blurring, which can be easily integrated. However, the non-differentiable nature of H.264 compression complicates training. To overcome this, we design an H.264 network inspired by DVMark [21] to mimic real compression, improving robustness.
3.4 Training Objectives
We design our objective function to balance the trade-off between visual quality and watermark decoding accuracy. We employ binary cross-entropy loss between the random message $m$ and the extracted message $\hat{m}$ from the generated video to ensure accurate watermark decoding:
$\mathcal{L}_{message}=-\frac{1}{N}\sum_{i=1}^{N}$ $m_{i}\cdot\log(\sigma(\hat{m}_{i}))$ (3)
$+(1-m_{i})\cdot\log(1-\sigma(\hat{m}_{i})),$
where $N$ is the total number of message bits, while $\sigma(\cdot)$ denotes the sigmoid function. In addition, to enhance the visual quality of the generated video, we choose the Watson-VGG perceptual loss [8], which is designed to avoid blurring effects and visual artifacts of generated images:
$\mathcal{L}_{vgg}=\text{VGG}(\mathcal{D}_{mod}({\mathcal{M}}(m),{\mathcal{E}}(v)),v),$ (4)
where $\mathcal{D}_{mod}$, $\mathcal{M}$, and $\mathcal{E}$ denote the modulated latent decoder, mapping network, and latent encoder, respectively, while $m$ and $v$ are the random message and input video.
Furthermore, we propose a weighted patch loss to address localized artifacts in the generated video. We first compute the mean absolute error (MAE) for video patches, resulting in a loss score map normalized between 0 and 1 using the softmax function. By multiplying the patch MAE by the score map, we create a weighted patch loss that places more emphasis on areas with higher MAE, where artifacts are likely to occur. The weighted patch loss is as follows:
$\mathcal{L}_{patch}$ $=\frac{1}{P}\sum_{i=1}^{P}|\hat{v}_{i}-v_{i}|\cdot\text{softmax}(|\hat{v}_{i}-v_{i}|),$ (5)
where $P$ denotes the number of patches. The final objective can be formulated as:
$\mathcal{L}=\lambda_{msg}\mathcal{L}_{msg}+\lambda_{vgg}\mathcal{L}_{vgg}+\lambda_{patch}\mathcal{L}_{patch},$ (6)
where $\lambda_{msg}$, $\lambda_{vgg}$, and $\lambda_{patch}$ are hyper-parameters that balance three loss functions.
Figure 3: Qualitative results with baselines. We show the visual quality of generated videos with baseline watermarking methods. The first row shows the crop of each image, while the second and third rows show the video frame itself and the difference map($\times$10) between the original and each method. Note that, unlike other approaches, VideoShield [12] embeds watermarks in the latent noise of the diffusion model, generating a video different from the original.
4 Experiments
4.1 Dataset
Training Dataset. We train our method on Panda-70M dataset [6], a large-scale collection of video frames recognized for its diversity across various content types. We randomly download 10K videos from this dataset. Following the approach of Stable Video Diffusion [3], we sample 8 frames from each video, with each frame resized to a resolution of $256\times 256$ for training.
Prompt Dataset. We evaluate text-to-video generation on the Vidprom dataset [34], using prompts generated by GPT-4 [1]. To evaluate our method, we randomly sample 100 prompts from the dataset, which do not contain personally identifiable information or offensive content.
4.2 Implementation Details
Video diffusion models are generally classified into two categories: DiT-based [26] and U-Net-based [27] architectures. To evaluate our method, we utilize Open-Sora [45] for the DiT architecture and DynamiCrafter [37] for the U-Net architecture, both of which are state-of-the-art open-source models, with Open-Sora excelling in text-to-video generation and DynamiCrafter in text-and-image-to-video generation. Our method is trained for over 40K iterations using the AdamW optimizer [20] on a single A100 GPU. We modulate 50% of the latent decoder layers to embed the message. For fair comparisons, we apply Bose–Chaudhuri–Hocquenghem (BCH) codes, a type of error correction code, to all baseline watermarking methods under the same setting to detect and correct minor bit errors. For more details, please refer to the supplementary material. The generated videos have a resolution of $512\times 320$ for DynamiCrafter and $640\times 640$ for Open-Sora, with both models generating 16 frames. For optimization, we set $\lambda_{msg}=0.8$, $\lambda_{vgg}=0.7$, $\lambda_{patch}=10$ in Eq. 6. Furthermore, we employ 100 random messages.
4.3 Evaluation
[…]
Baselines. To the best of our knowledge, there is no watermarking method for video diffusion models that preserves the original content. Therefore, for a fair comparison, we evaluate seven methods across three types of watermarking methods: 1) HiDDeN [46] and WAM [30] are watermarking methods for images. 2) Blind [23] and REVMark [44] are watermarking methods for videos. 3) Stable Signature [10], WOUAF [16], and VideoShield [12] are watermarking methods for diffusion models. For both Stable Signature and WOUAF, we train Open-Sora [45] and DynamiCrafter [37] using the same training set as our method. Additionally, we decode the message from each video frame for image-level watermarking methods [46, 30, 10, 16].
4.4 Results
Video Quality and Bit Accuracy. There is a trade-off between video quality and bit accuracy, making it a key challenge in digital watermarking to improve both simultaneously. Consequently, for a fair comparison, we measure the video quality when bit accuracy is maintained close to 100%. We first compare the generated video quality and bit accuracy of our method against other approaches, as shown in Figure 3 and Table 1. We observe that image-based watermarking methods [46, 30, 10, 16] do not fully address temporal consistency, as they independently embed the invisible message into each frame of the generated video. The performance drop of image diffusion model watermarking methods [10, 16] is particularly pronounced when applied to video diffusion models, as they are designed for image-based models that ignore temporal information. Video watermarking methods [23, 44] maintain relatively high bit accuracy and video quality but exhibit inferior performance compared to our method. Our method preserves video quality, including temporal consistency, while achieving the highest bit accuracy (more details in Section 4.5).
Watermark Capacity. In this section, we analyze the video quality and bit accuracy based on the length of the embedded message. A higher watermark bit capacity enables the registration and tracking of more users’ copyright information, enhancing the practicality of watermarking methods. We compare our method with others using 100 different messages for 32-bit and 96-bit capacities. Furthermore, we embed a 512-bit message into our method, demonstrating its ability to maintain both video quality and bit accuracy at higher capacities. As shown in Table 1, all methods exhibit a decline in bit accuracy and visual quality as the message length increases. Notably, our method achieves superior video quality, even when embedding a 512-bit message, outperforming other methods with a 32-bit capacity. Therefore, experiments with higher capacities for other methods are unnecessary. This highlights the superiority of our method in embedding high-capacity watermarks while preserving video quality better than other approaches.
Open Sora DynamiCrafter
Method Capacity (Bit) Bits Acc.(%) $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ tLP $\downarrow$ FVD $\downarrow$ Bits Acc.(%) $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ tLP $\downarrow$ FVD $\downarrow$
HiDDeN [46] 32 99.78 30.19 0.877 0.237 5.310 2980.7 99.82 30.83 0.888 0.178 7.596 1947.7
WAM [30] 32 99.99 30.84 0.917 0.170 2.738 435.4 99.99 30.92 0.911 0.156 3.777 317.1
Blind [23] 32 99.71 29.88 0.888 0.181 2.615 363.2 99.63 29.45 0.882 0.179 2.348 389.5
96 99.12 28.32 0.842 0.194 2.937 401.2 99.05 28.05 0.854 0.187 3.019 416.3
REVMark [44] 32 99.99 29.26 0.905 0.128 0.317 133.6 99.99 29.02 0.913 0.113 0.548 130.9
96 99.98 29.24 0.891 0.134 0.497 249.1 99.99 28.85 0.898 0.114 0.631 189.1
Stable Signature [10] 32 99.96 27.33 0.821 0.211 3.387 1094.8 99.02 26.78 0.797 0.223 4.353 954.8
96 95.88 26.46 0.803 0.259 4.019 1688.5 95.68 26.32 0.788 0.240 4.603 1262.8
WOUAF [16] 32 99.28 28.62 0.873 0.166 4.856 993.7 99.62 28.38 0.881 0.125 5.763 1469.4
96 96.25 28.13 0.858 0.177 5.053 1196.2 96.22 28.23 0.877 0.130 5.989 1523.2
VideoShield [12] 512 99.52 - - - - - 99.40 - - - - -
LVMark (Ours) 32 99.99 32.65 0.938 0.094 0.167 63.70 99.99 32.22 0.932 0.058 0.141 43.27
96 99.96 32.60 0.935 0.099 0.198 70.46 99.99 32.06 0.925 0.063 0.169 46.83
512 99.81 31.74 0.918 0.105 0.229 83.98 99.98 31.63 0.914 0.066 0.189 68.94
Table 1: Quantitative results. We present the quantitative results for Open-Sora [45] and DynamiCrafter [37] with various watermarking methods: image watermarking methods [46, 30], video watermarking methods [23, 44], diffusion model watermarking methods [10, 16, 12]. The evaluation includes image metrics: PSNR, SSIM, LPIPS, video metrics: tLP [7], FVD [33], and bit accuracy on 32, 96 and 512-bit messages. Note that since VideoShield [12] generates a different video from the original, the visual quality metrics in the table are not meaningful and are left blank. Best performance in bold, second in underline, lower in gray.
Capacity Image Attack Video Attack Model Attack
(Bit) None
Crop
(50%)
Rotation
($+\pi/6$)
Blur
(std = 1.0)
JPEG
(50%)
Frame Swap
(p=0.5)
Frame Avg.
(n=6)
Frame Drop
(p=0.5)
H.264
(crf=24)
Combine
Gaussian / (PSNR)
(p=0.2)
Prune / (PSNR)
(p=0.1)
Retrain / (PSNR)
Open Sora HiDDeN [46] 32 99.10 99.07 50.24 63.89 59.47 99.10 98.93 99.75 55.66 55.03 - - -
WAM [30] 32 99.99 99.94 53.74 99.95 91.94 99.99 99.97 99.99 87.60 58.03 - - -
Blind [23] 96 99.12 88.51 87.32 69.24 67.53 89.51 59.53 90.12 84.13 88.19 - - -
REVMark [44] 96 99.98 82.91 59.47 99.59 94.65 99.74 99.98 98.09 93.49 81.20 - - -
Stable Signature [10] 96 95.88 92.49 94.15 89.80 83.61 95.88 93.29 94.14 82.63 76.32 93.15 / (24.67) 93.99 / (20.03) 50.19 / (26.51)
WOUAF [16] 96 96.25 88.51 95.33 87.80 87.75 96.25 94.56 93.26 90.15 86.32 93.68 / (26.45) 92.93 / (19.67) 85.47 / (28.15)
VideoShield [12] 512 99.52 96.93 52.89 99.49 98.12 99.34 99.64 93.16 97.92 75.25 96.32 97.19 -
LVMark (Ours) 32 99.99 99.96 99.95 99.96 99.84 99.99 99.98 99.99 99.84 99.83 99.52 / (29.02) 99.85 / (22.72) 99.16 / (32.69)
96 99.96 99.91 99.82 99.95 99.77 99.96 99.96 99.96 99.76 99.81 99.46 / (28.97) 99.84 / (22.70) 98.99 / (32.58)
512 99.81 99.74 99.51 99.66 99.19 99.81 99.04 99.24 98.20 99.19 99.29 / (28.88) 99.81 / (22.53) 97.67 / (31.72)
DynamiCrafter HiDDeN [46] 32 99.02 98.82 50.38 62.64 58.81 99.02 98.91 99.81 56.01 55.82 - - -
WAM [30] 32 99.99 99.78 63.11 99.98 94.34 99.99 99.98 99.99 85.27 64.20 - - -
Blind [23] 96 99.05 84.26 85.75 69.29 61.48 85.19 57.25 86.24 85.63 85.02 - - -
REVMark [44] 96 99.99 81.95 58.27 99.88 95.86 99.79 99.99 97.83 94.52 80.29 - - -
Stable Signature [10] 96 95.68 92.11 93.86 89.03 84.11 95.68 93.23 94.01 81.17 75.32 93.68 / (23.92) 94.15 / (19.94) 50.24 / (26.35)
WOUAF [16] 96 96.22 86.25 94.84 85.80 87.47 96.22 94.42 93.10 88.90 85.09 94.13 / (25.11) 94.07 / (18.96) 83.43 / (28.21)
VideoShield [12] 512 99.40 96.65 50.11 99.20 97.99 99.26 99.35 92.57 97.60 71.96 96.05 96.81 -
LVMark (Ours) 32 99.99 99.98 99.98 99.98 99.98 99.99 99.99 99.99 99.98 99.74 99.74 / (26.45) 99.90 / (21.24) 97.36 / (32.25)
96 99.99 99.98 99.98 99.98 99.95 99.99 99.97 99.99 99.94 99.50 99.72 / (26.41) 99.87 / (21.19) 97.24 / (32.02)
512 99.98 99.90 99.93 99.74 99.94 99.98 99.04 99.36 99.75 99.16 99.62 / (26.26) 99.76 / (21.03) 95.73 / (31.59)
Table 2: Robustness comparisons. We present bit accuracy of watermarking methods under various attacks. Image and video methods [46, 30, 23, 44] are not applicable to model attacks, and VideoShield [12] does not fine-tune the diffusion model, leaving them blank.
Robustness for Video Distortions. A key priority in digital watermarking is robustness against distortions. In this section, we evaluate our method’s robustness to distortions in videos. Table 2 shows the bit accuracies of our method and baselines against various distortions. Leveraging spatio-temporal information in the 3D wavelet domain, our watermark decoder achieves high bit accuracy across all evaluated attacks. Notably, it demonstrates exceptional robustness against compressions such as H.264, by extracting messages from low-frequency components. We note that since all methods except video watermarking methods [23, 44] are applied independently per frame, it is naturally robust against frame swap.
Robustness for Model Distortions. To mitigate the risk of unauthorized model usage, we assess the robustness of our method against model distortions. Since strong Gaussian noise or layer pruning degrades video quality, unauthorized users are unlikely to apply significant distortions. Maintaining high bit accuracy despite quality degradation demonstrates the robustness of the watermark decoder. Therefore, we evaluate bit accuracy under compromised video quality. Additionally, we evaluate bit accuracy when an unauthorized user retrains the model using a specific watermark key. The right side of Table 2 reports bit accuracy and PSNR under model attacks. Our method ensures the presence of messages even when the diffusion model is distorted. Model distortion alters parameters, causing variations in the generated video details. By extracting watermarks from both the wavelet and RGB domains, and training with varying watermark keys and distortion layers, our decoder becomes resilient to input variations, ensuring robustness against model distortion.
4.5 Ablation Study
For the ablation study, we follow the same setup as in Section 4.2 with 512-bit messages. For additional experimental results, please refer to the supplementary materials.
Figure 4: The impact of weight modulation rate. Each point represents the metrics for different modulation rates: 0%, 25%, 50%, 75%, and 100%.
Importance-based Weight Modulation. As mentioned in Section 3.2, we selectively modulate the top $k$ of layers with the least impact on video quality to embed messages while minimizing visual degradation. As shown in Figure 4, as the modulation rate $k$ approaches 100%, bit accuracy improves at the cost of video quality, whereas setting $k$ as 0% results in failed bit extraction, illustrating a trade-off. We find an optimal balance between bit accuracy and video quality when $k$ is set to 50.
DWT Type Bits Acc. $\uparrow$ PSNR $\uparrow$ tLP $\downarrow$ FVD $\downarrow$
Open Sora 3D DCT 98.74 30.73 0.249 104.8
3D FFT 97.98 31.28 0.232 85.37
2D DWT 96.89 30.97 0.410 175.4
3D DWT (Ours) 99.81 31.74 0.229 83.98
DynamiCrafter 3D DCT 99.21 30.72 0.293 114.7
3D FFT 95.86 31.42 0.197 72.35
2D DWT 96.63 30.89 0.324 143.9
3D DWT (Ours) 99.98 31.63 0.189 68.94
Table 3: Comparison of different frequency types. Evaluation of bit accuracy, PSNR, tLP, and FVD across various frequency transformation techniques.
Bit Accuracy (%) $\uparrow$
Domain Type PSNR $\uparrow$ None H264 (crf24)
Open Sora RGB only 30.87 94.87 86.99
High Freq only 33.74 50.62 50.54
Low Freq only 33.85 50.23 50.13
High Freq + RGB 30.50 95.46 89.40
Low Freq + RGB (Ours) 31.74 99.81 98.20
DynamiCrafter RGB only 30.49 94.31 85.90
High Freq only 33.19 51.29 51.21
Low Freq only 33.23 50.93 51.78
High Freq + RGB 30.35 94.78 88.93
Low Freq + RGB (Ours) 31.63 99.98 99.75
Table 4: Comparison of different domain types. Evaluation of PSNR and bit accuracy using different frequency components.
Watermark Decoder with 3D DWT. Temporal consistency is crucial in video watermarking and generation. To enhance this, we use the 3D Wavelet Transform (3D DWT), which captures spatio-temporal features by decomposing videos into wavelet representations. Table 3 compares 3D Discrete Cosine Transform, 3D Fast Fourier Transform, 3D DWT, and 2D DWT. Our experiments show that 3D DWT outperforms other methods in bit accuracy and video quality due to its superior localization in time and frequency. In contrast, 2D DWT, which only captures spatial information, results in a loss of temporal consistency, degrading video quality metrics such as tLP and FVD. Based on these findings, we utilize 3D DWT in our watermark decoder.
Watermark Decoder with Low-frequency. As mentioned in Section 3.1, we utilize only the low-frequency subband ($LLL$) among the 3D wavelet subbands. While this achieves high PSNR, relying solely on the frequency domain limits bit accuracy, as shown in Table 4. By combining RGB features with the low-frequency subband ($LLL$), we significantly enhance bit accuracy without compromising PSNR. Our method is particularly robust against compression techniques like H.264, which preserve low-frequencies.
Figure 5: Visual impact of weighted patch loss. We visualize a video frame trained with and without weighted patch loss. To enhance clarity, we crop the local regions containing artifacts. Both results used VGG Loss [8] with the default setting.
Impact of Weighted Patch Loss. As described in Section 3.4, weighted patch loss mitigates localized artifacts, enhancing overall video quality. VGG Loss [8] is often used to improve the perceptual quality of images, but it has limitations in enhancing PSNR, which measures pixel-wise differences. To address this, combining VGG Loss with Mean Absolute Error (MAE) Loss, which focuses on pixel-wise differences, can be effective. However, as shown in Figure 5, our experiments reveal that using both losses together introduces localized artifacts, as the embedded watermark in the video creates artifacts that degrade visual quality. Table 5 demonstrates that weighted patch loss enables higher bit accuracy while preserving video quality, outperforming MAE Loss. This highlights the effectiveness of weighted Patch Loss, considering the trade-off between bit accuracy and visual quality.
Open Sora DynamiCrafter
Bits Acc. $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ Bits Acc. $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$
VGG Loss 99.78 30.86 0.107 99.86 30.74 0.071
VGG + MAE Loss 99.80 31.12 0.139 99.82 31.03 0.107
VGG + Patch Loss (Ours) 99.81 31.74 0.105 99.98 31.63 0.066
Table 5: Ablation on weighted patch loss. Evaluation of bit accuracy and visual quality across loss functions.
5 Conclusion
We proposed LVMark, a robust watermarking method for video diffusion models. Our method captures spatio-temporal information by effectively fusing the 3D wavelet domain and the RGB spatial domain for accurate message extraction from generated videos, even in the case of video and model modifications. Additionally, our method generates high-quality watermarked videos by selectively modulating latent decoder weights according to the importance of each layer. By balancing the trade-off between video quality and bit accuracy with tailored objective functions, including weighted patch loss, LVMark successfully embeds watermarks into video diffusion models. Experimental results show that our method is the first practical video diffusion watermarking approach, achieving a high capacity of 512-bit while preserving original content and ensuring robust watermark retention against various attacks, including model-based ones.
Limitations & Future Work. Our method excels in watermarking video diffusion models. However, training requires significant memory, around 25GB, due to the large size of the diffusion models and videos. Future work could explore memory-efficient watermarking techniques.